# Cardiac Copilot: Automatic Probe Guidance for Echocardiography with World Model

### v1

@@ 1 Introduction
With technological advancements [5, 8, 10, 11, 12, 19, 20], AI-driven systems show immense potential to significantly improve the efficiency of medical ultrasound processes. For example, Droste et al. [1] proposed an US-GuideNet to provide only rotation movement guidance signal for acquiring standard planes in routine freehand obstetric ultrasound scanning. However, in routine ultrasound scanning, the first step often involves moving close to the target plane through translation. Thus, providing a six-dimension (6D) guidance signal is essential, which includes signals for both translation and rotation. Narang et al. [13] proposed an AI-driven commercial solution to guide nurses to acquire echocardiogram by suggesting approximate adjustment directions in one degree of freedom. They further conducted clinical trials and found that it could assist nurses in locating cardiac target planes that meet diagnostic requirements. Nonetheless, the specifics of the methodological implementation remain undisclosed, impeding progress in cardiac ultrasound domain. Later, Shida et al. [15, 16] investigated the potential of using manually predefined search strategies to acquire the Parasternal Long-Axis (PLAX) plane. The authors conducted a small-scale experiment on five healthy subjects, with the upper limit of the success rate being only 63.3%. This indicates that the rule-based guidance strategy lacks generalizability and practicality. Despite on-going efforts, there is still no publicly available research on cardiac probe navigation, which hampers further advancement in the field.
@@ 2.1 Target-oriented Guidance Framework
Figure 2: Framework diagram of target-oriented guidance framework. Left is the policy network which provides the basic guidance signal for locating the target plane. Right is the Cardiac Dreamer which foresees states reached by executing actions output by the policy network and refines the actions based on these states. The details of the action combining operator is shown in Eq. 8.
@@ 2.2 Foresee States by Cardiac Dreamer
Training. Given the processed demonstration dataset $D_{i}^{{}^{\prime}}$, we create the permutation dataset $D^{{}^{\prime\prime}}$ with $d_{i}^{{}^{\prime\prime}}=\{(\mathbf{S}_{t_{1}},\mathbf{S}_{t_{2}},\mathbf{\hat{a}}_{t_{1}\to t_{2}},\mathbf{\hat{a}}_{t_{1}},\mathbf{\hat{a}}_{t_{2}})|t_{1}\neq t_{2}\}$, where $\mathbf{\hat{a}}_{t_{1}\to t_{2}}$ is relative positional between $S_{t_{1}}$ and $S_{t_{2}}$, and $\mathbf{\hat{a}}_{t_{1}}$, $\mathbf{\hat{a}}_{t_{2}}$ is relative positional between $S_{t_{1}}$, $S_{t_{2}}$ and the target plane. Then, $\mathbf{S}_{t_{1}}$ and $\mathbf{S}_{t_{2}}$ are processed by the feature encoder $E(\cdot)$ inside the policy network to yield $\mathbf{f}_{t_{1}}$ and $\mathbf{f}_{t_{2}}$:
$\mathbf{f}_{t_{1}}=E(\mathbf{S}_{t_{1}}),\ \ \mathbf{f}_{t_{2}}=E(\mathbf{S}_{t_{2}}).$ (3)
@@
The generated feature $\mathbf{f}_{t_{2}}^{{}^{\prime}}$ is supervised by $\mathbf{f}_{t_{2}}$ with mean squared loss. Through this method, we aim for Cardiac Dreamer to learn the spatial relationships between various planes, thereby gaining a comprehensive understanding of cardiac structures. Furthermore, the generated $\mathbf{f}_{t_{2}}^{{}^{\prime}}$ is used to predict the guidance signal $\mathbf{a}_{t_{2}}^{{}^{\prime}}$ for locating the target plane:
$\mathbf{a}_{t_{2}}^{{}^{\prime}}=G(\mathbf{f}_{t_{2}}^{{}^{\prime}}).$ (5)
@@
Inference. As shown in Fig. 2, the Cardiac Dreamer can be utilized to foresee states reached by executing actions output by the policy network and to refine the actions based on these states. Given the initial action $\mathbf{a}_{t,0}$ generated by the policy network at time $t$ and the action $\Delta\mathbf{a}_{t,0}$ predicted based on the feature $\mathbf{f}_{t,0}$ generated by Cardiac Dreamer, a more precise guidance signal $\mathbf{a}_{t,1}$ for reaching the target plane can be obtained by combining $\Delta\mathbf{a}_{t,0}$ onto $\mathbf{a}_{t,0}$:
$\Delta\mathbf{a}_{t,0}=G(I(\mathbf{f}_{t},\mathbf{a}_{t,0})),\ \ \mathbf{a}_{t,1}=U^{-1}(U(\mathbf{a}_{t,0})U(\Delta\mathbf{a}_{t,0})),$ (8)
@@ 3.1 Setup
In the experimental setup, three standard echocardiographic planes, i.e., the Parasternal Long-Axis (PLAX), the Parasternal Short-Axis Aortic Valve (PSAX-AV), and the Parasternal Short-Axis Mitral Valve (PSAX-MV), were chosen for their clinical importance, offering critical insights into cardiac anatomy and functionality. The ultrasound image data was acquired from health adult male volunteers using a Franka Panda robot arm with an ultrasound probe (M5S probe, GE Vivid E7 machine) manipulated by professional echocardiographers, meanwhile the pose information of the probe was also recorded synchronously by utilizing Robot Operation System (ROS). Overall, approximately 188K cardiac ultrasound images paired with probe motion were collected from 125 routine clinical scans, forming a diverse dataset for training the proposed Cardiac Copilot models. The whole data collecting process was approved and supervised by The Tsinghua University Science and Technology Ethics Committe. Data processing involved aligning relative pose coordinates of each image with target planes which were identified by professionals, ensuring a unified coordinate system for each plane. The dataset was split into 110 scans (about 151K images) for training and 15 scans (about 37K images) for testing.
@@ 3.3 Evaluation Metrics
Figure 3: Detailed analysis. The change of predicted absolute error of different samples with respect to the mean pose difference to target plane on three standard planes. The dashed line represents the mean value, and the shading represents the standard deviation.

### v2

@@ 1 Introduction
With technological advancements [5, 8, 10, 11, 12, 19], AI-driven systems show immense potential to significantly improve the efficiency of medical ultrasound processes. For example, Droste et al. [1] proposed an US-GuideNet to provide only rotation movement guidance signal for acquiring standard planes in routine freehand obstetric ultrasound scanning. However, in routine ultrasound scanning, the first step often involves moving close to the target plane through translation. Thus, providing a six-dimension (6D) guidance signal is essential, which includes signals for both translation and rotation. Narang et al. [13] proposed an AI-driven commercial solution to guide nurses to acquire echocardiogram by suggesting approximate adjustment directions in one degree of freedom. They further conducted clinical trials and found that it could assist nurses in locating cardiac target planes that meet diagnostic requirements. Nonetheless, the specifics of the methodological implementation remain undisclosed, impeding progress in cardiac ultrasound domain. Later, Shida et al. [15, 16] investigated the potential of using manually predefined search strategies to acquire the Parasternal Long-Axis (PLAX) plane. The authors conducted a small-scale experiment on five healthy subjects, with the upper limit of the success rate being only 63.3%. This indicates that the rule-based guidance strategy lacks generalizability and practicality. Despite on-going efforts, there is still no publicly available research on cardiac probe navigation, which hampers further advancement in the field.
@@ 2.1 Target-oriented Guidance Framework
Figure 2: Diagram of target-oriented guidance framework. Left is the policy network which provides the basic guidance signal for locating the target plane. Right is the Cardiac Dreamer which foresees states reached by executing actions output by the policy network and refines the actions based on these states. The details of the action combining operator is shown in Eq. 6.
@@ 2.2 Foresee States by Cardiac Dreamer
Given the processed demonstration dataset $D_{i}^{{}^{\prime}}$, we create the permutation dataset $D^{{}^{\prime\prime}}$ with $d_{i}^{{}^{\prime\prime}}=\{(\mathbf{S}_{t_{1}},\mathbf{S}_{t_{2}},\mathbf{\hat{a}}_{t_{1}\to t_{2}},\mathbf{\hat{a}}_{t_{1}},\mathbf{\hat{a}}_{t_{2}})|t_{1}\neq t_{2}\}$, where $\mathbf{\hat{a}}_{t_{1}\to t_{2}}$ is relative positional between $S_{t_{1}}$ and $S_{t_{2}}$, and $\mathbf{\hat{a}}_{t_{1}}$, $\mathbf{\hat{a}}_{t_{2}}$ is relative positional between $S_{t_{1}}$, $S_{t_{2}}$ and the target plane. Then, $\mathbf{S}_{t_{1}}$ is processed by the feature encoder $E(\cdot)$ inside the policy network to yield $\mathbf{f}_{t_{1}}$:
$\mathbf{f}_{t_{1}}=E(\mathbf{S}_{t_{1}}).$ (3)
@@
We aim for Cardiac Dreamer to learn the spatial relationships between various planes, thereby gaining a comprehensive understanding of cardiac structures. Furthermore, the generated $\mathbf{f}_{t_{2}}^{{}^{\prime}}$ is used to predict the guidance signal $\mathbf{a}_{t_{2}}^{{}^{\prime}}$ for locating the target plane:
$\mathbf{a}_{t_{2}}^{{}^{\prime}}=G(\mathbf{f}_{t_{2}}^{{}^{\prime}},\mathbf{q}_{i}).$ (5)
@@ 3.1 Setup
In the experimental setup, three standard echocardiographic planes, i.e., the Parasternal Long-Axis (PLAX), the Parasternal Short-Axis Aortic Valve (PSAX-AV), and the Parasternal Short-Axis Mitral Valve (PSAX-MV), were chosen for their clinical importance, offering critical insights into cardiac anatomy and functionality. The ultrasound image data was acquired from health adult male volunteers using a Franka Panda robot arm with an ultrasound probe (M5S probe, GE Vivid E7 machine) manipulated by professional echocardiographers, meanwhile the pose information of the probe was also recorded synchronously by utilizing Robot Operation System (ROS). Overall, approximately 188K cardiac ultrasound images paired with probe motion were collected from 125 routine clinical scans, forming a diverse dataset for training the proposed Cardiac Copilot models. The whole data collecting process was approved and supervised by The Tsinghua University Science and Technology Ethics Committe. Data processing involved aligning relative pose coordinates of each image with target planes which were identified by professionals, ensuring a unified coordinate system for each plane. The dataset was split into 110 scans (about 151K images) for training and 15 scans (about 37K images) for testing. Note that the test set included individuals not seen during training to evaluate the algorithm’s generalization.
@@ 3.3 Evaluation Metrics
Figure 3: Detailed analysis. For optimal viewing, please consider enlarging and viewing in color. The change of predicted absolute error of different samples with respect to the mean pose difference to target plane on three standard planes. The ’mean pose difference’ refers to calculating the mean absolute error between the pose of the current plane and that of the target plane. The dashed line represents the mean value, and the shading represents the standard deviation.